The answer key, also known as Tricky and Twisted Math has been reverted to draft because Kazuma, Annis, and I are doing a class thing and we need the students to figure it out on their own rather than cheat so, sorry!
Showing posts with label Math Problem of the Day. Show all posts
Showing posts with label Math Problem of the Day. Show all posts
5/23/14
Goodbye Answers
Labels:
Annis Jeffrey,
Kazuma,
Math (duh),
Math Problem of the Day,
pages,
scheduling,
Strix Spell,
The Laws of Not Pi
5/22/14
Problem of the Day #10
Hooray, we are on 10!
So, more patterns, oh, and this has something to do with twin primes.
What comes next in this sequence? 4, 6, 12, 18, 30, 42, 60, 72, 102, 108,...?
Have fun solving!
Problem of the Day #9
Sorry for not posting yesterday.
There were some complications but, I'm here today!
I like patterns so here is another riddle for patterns.
1 = 3, 2 = 3, 3 = 5, 4 = 4, 5 = 4, 6 = 3, 7 = 5, 8 = 5, 9 = 4, 10 = 3, 11 = ?, 12 = ?
What do eleven and twelve equal?
There were some complications but, I'm here today!
I like patterns so here is another riddle for patterns.
1 = 3, 2 = 3, 3 = 5, 4 = 4, 5 = 4, 6 = 3, 7 = 5, 8 = 5, 9 = 4, 10 = 3, 11 = ?, 12 = ?
What do eleven and twelve equal?
5/20/14
Problem of the Day #8
I'm covering for Kazuma because we usually work together on posts but I was late for the meeting. It is completely my fault.
Figure out the next part of this pyramid:
1
11
21
1211
111221
312211
13112221
So, what comes next? This will help you sharpen your pattern senses.
5/19/14
Problem of the Day #5
Ok, I forgot to cover for Kazuma so I'm doing three problems today. Kazuma will be helping me.
The problem:
If one nickel is worth 5 cents, how much is half of one half of a nickel worth?
If one nickel is worth 5 cents, how much is half of one half of a nickel worth?
5/16/14
Problem of the Day #4
Chiyo is picking between 7 different candy canes to put on her Christmas tree. However, she only wants 3 to go on the tree. What is the number of combinations of candy canes Chiyo can pick if she is only picking 3 candy canes at a time?
5/15/14
Problem of the Day #3
The cylinder's diameter was 4 inches and it was 8 inches tall. Shiro and Jamie were trying to find the volume of the cylinder Shiro says that you need to do 3.14 times 2 squared times 8. Jamie says that you need to do 3.14 times 4 squared times 8. Who's right? Explain.
Comment your answers!
Comment your answers!
5/14/14
Problems of the Day #2
Jake bought two rectangular containers of coffee powder. The decaf coffee container is 1/4 feet long, 2 inches wide, and 7 inches tall. The regular coffee container is 3 inches long, 1/4 feet wide, and 3/4 feet tall. Which coffee container would hold more coffee powder?
Comment your answers!
Comment your answers!
5/13/14
The History of Probability
Probability: the measure of the likelihood that an event will occur.
There are two types of events in probability: simple events and compound events. Simple events... are simple. Simple events are the likelihood of only one event. A simple event would be like: What is the probability of rolling a 4 on a standardized die? That would be 1/6. If you don't get that, to calculate simple events, create a ratio of the number of times your number shows up (4 only appears once on the die)to the number of possible outcomes(in this case 6, there are 6 faces on a die). Anyway, to the history. Probability was invented by two French mathematicians, Blaise Pascal and Pierre de Fermat. Pascal should technically be credited more because he really came up with the idea. See, Pascal was prompted with the idea of probability by a popular dice game where you would throw a pair of dice 24 times. You needed to bet on whether or not the dice would both land on 6 at least once. Pascal then began to correspond(exchange letters) with Fermat. This eventually formulated the basis for theoretical probability. Later, more mathematicians wrote books and developed different styles of probability, so really, Pascal and Fermat's creation was a simple event. #badpun... Anyways, compound events are events with at least two events, like calculating the probability of a 5 on a die and red on a spinner. There are 3 types of compound events: mutually exclusive events, independent events, and dependent events. Mutually exclusive events are two events that cannot both occur. The formula for that is P(A or B)= P(A)+P(B). For example if I wanted to find the probability of landing on blue or red on a blue, red, yellow, and green spinner, I would solve it like this. P of blue=1/4. P of red= 1/4. 1/4 +1/4=1/2. Therefore, the probability of landing on a blue or red is 1/2. A type of mutually exclusive event is a complementary event. Complementary events are mutually exclusive events that probabilities together are always equal to 1. 1 in the probability world means that it is certain to happen. Independent events are two events that do not affect each other. The formula for that is P(A and B)= P(A) x P(B). Basically, when calculating the probability of a die and spinner, you are calculating independent events. Like the probability of 2 and green. Dependent events are two events that has one event depending on the other. The formula for that is P(A and B)= P(A) x P(B)given A. For example, what is the probability of picking two pink marbles in a row if there are 7 pink and 3 orange marbles in a bowl. Remember, the odds of getting a pink change when you have already picked one out. Answer the marble probability question in the comments below. See Ya!
Problem of the Day #1
Kitti bought 3 water bottles, 2 crayon boxes, and 4 strawberry Japanese sodas. The water bottles are half the cost of the soda each and the crayon boxes are 3/4 the cost of the bottles each. About how much did Kitti spend if the sodas are $1.50 each?
Comment your answers!
Comment your answers!
Subscribe to:
Posts (Atom)




